# Identification of Semiparametric Panel Multinomial

### v2

@@ 4 General Econometric Framework of Multi-Index Single-Crossing Conditions
Similar to Example 2, define $\xi_{i}:=W_{i1}-Z_{i}^{{}^{\prime}}\gamma_{0}$ as the residual from the reduced-form linear projection of $W_{i1}$ on the instrumental variables $Z_{i}$, and sssume that the endogeneity between $\epsilon_{i}$ and $W_{i1}$ is captured by the control function $\text{med}\left(\left.\epsilon_{i}\right|W_{i},Z_{i}\right)=\text{med}\left(\left.\epsilon_{i}\right|Z_{i},\xi_{i}\right)=\lambda\left(\alpha_{0}\xi_{i}\right)$, where $\lambda$ being an increasing function with normalization $\lambda\left(0\right)=0$. Writing $\overline{Z}_{i}:=\left(W_{i1},Z_{i}\right)$ and $\overline{\gamma}_{0}:=\left(\alpha_{0},-\alpha_{0}\gamma_{0}\right)^{{}^{\prime}}$, we have
$W_{i}^{{}^{\prime}}\beta_{0}>0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}>0\quad\Rightarrow\quad\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ $>0.$
$W_{i}^{{}^{\prime}}\beta_{0}\leq 0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}\leq 0\quad\Rightarrow\quad\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ $=0.$
which can be viewed as a MISC condition with a weak $\leq$” side and $h_{0}\left(X_{i}\right):=\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ given by the conditional median instead of conditional expectation in Example 2.
Given the MISC conditions (15), taking the logical contraposition again yields identifying restrictions, which can be encoded algebrically in a similar manner as in (6) and (9). Specifically, let $G$ again be a ono-sided sign-preserving function as in (9), and define
$\lambda_{j}\left(W_{i};\theta\right):=\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{W_{ij}^{{}^{\prime}}\theta_{j}\leq 0\right\}.$
@@
$h_{0}\left(X_{i}\right)>0$ $\ \Rightarrow\ \text{NOT}\ \left\{W_{ij}^{{}^{\prime}}\theta_{j}\geq 0\ \forall j\right\}.$
Furthermore, with $Q^{G}:=Q_{+}^{G}+Q_{-}^{G}$ and
$Q_{+}^{G}\left(\beta\right)$ $:=\mathbb{E}\left[G\left(h_{0}\left(X_{i}\right)\right)\lambda_{j}\left(W_{i};\beta\right)\right],$
$Q_{-}^{G}\left(\beta\right)$ $:=\mathbb{E}\left[G\left(-h_{0}\left(X_{i}\right)\right)\lambda_{j}\left(-W_{i};\beta\right)\right],$ (16)
we have $Q^{G}\left(\beta\right)\geq Q^{G}\left(\beta_{0}\right)=0$.
Proposition 2 generalizes our key identification result (Theorem 1). Notice that Proposition 2 applies to all functionals $h_{0}$ on the conditional distribution of $y_{i}$ given ${\bf X}_{i}$ that satisfy the MISC conditions.
One could also proceed with the two-step estimation procedure described in Section 3. Given a first-stage nonparametric estimator $\hat{h}$ of $h_{0}$, we can estimate $\beta_{0}$ (or the identified set) based on the sample criterion $\hat{Q}^{G}:=\hat{Q}_{+}^{G}+\hat{Q}_{-}^{G}$ with
$\hat{Q}_{+}^{G}\left(\beta\right):=\frac{1}{n}\sum_{i=1}^{n}G\left(\hat{h}\left(X_{i}\right)\right)\lambda_{j}\left(W_{i};\beta\right)$
and $\hat{Q}_{-}^{G}$ similarly defined. Note again, that the maximum score estimator, which only applies to the single-index setting, cannot be similarly generalized in this manner.
@@ 5 Simulation
We now switch back to the panel multinomial choice model introduced in Section 2, which is our main model of interest in this paper, and examine the finite sample performance of our proposed estimator. For each DGP, we run $M=1,000$ simulations of model (1) with the following utility specification:
@@
where $A_{i0}$ is an unobserved scale fixed effect that captures agent-level heteroskedasticity in utilities, and $A_{ij}$ is an unobserved location shifter specific to each agent-product pair. The ability to deal with nonlinear dependence caused by the unobservable fixed effects in a robust way differentiates our method from others. To allow for such dependence, we generate correlation between the observable characteristics ${\bf X}_{i}$ and the fixed effects ${\bf A}_{i}$ via a latent variable $Z$ [^10] We draw $Z_{i}\sim\mathcal{N}$$\left(0,1\right)$ and let $A_{i2}=\left[Z_{i}\right]_{+}$. We construct $X_{ijt,2}=W_{ijt}+Z_{i}$ with $W_{ijt}\sim\mathcal{N}\left(0,2J\right)$. Thus, $X$ and $A$ are correlated via $Z$. The DGPs for the rest of ${\bf A}$ and ${\bf X}$ are: $A_{i0}\sim\mathcal{U}\left[2,2.5\right]$, $A_{i1}\equiv 0$, $A_{ij}\sim\mathcal{U}\left[-0.25,0.25\right]$ for $j\geq 3$, $X_{ijt,1}\sim\mathcal{U}\left[-1,1\right]$, $X_{ijt,d}\sim\mathcal{N}\left(0,1\right)$ for $d\geq 3$. Furthermore, we set $\overline{\beta}_{0}=\left(2,1,...,1\right)^{{}^{\prime}}\in\mathbb{R}^{D}$ and $\beta_{0}=\overline{\beta}_{0}/\left\|\overline{\beta}_{0}\right\|$, and draw $\epsilon_{ijt}\sim TIEV\left(0,1\right)$. To summarize, for each of the $M=1,000$ simulations we first generate $\left(\beta_{0},{\bf X}_{it},{\bf A}_{i},\boldsymbol{\epsilon}_{it}\right)$ for all $it$ combinations. Then we calculate the individual choice ${\bf Y}$ matrix according to model (1). Next, we compute $\hat{\beta}$ from the simulated observable data of $\left({\bf X},{\bf Y}\right)$, where for the first stage we use nonparametric regression with second-order polynomial basis functions with $\ell_{1}$-regularization and 10-fold cross validation to estimate $\gamma$ (thus $G\left(\gamma\right)$), and for the second stage we apply the adaptive-grid algorithm detailed in Section 3.2 to find $\hat{\beta}$. Finally, we evaluate the performance of $\hat{\beta}$ against the true $\beta_{0}$.  [^11] In Appendix D, we provide additional simulation results. Specifically, we first present a graphical illustration of the identified set $B_{0}$ based on the population criterion (10). Second, we inspect how our estimator performs without point identification. Lastly, we vary $\left(D,J,T\right)$ to examine how robust our method is against various simulation specifications.
@@
For the baseline configuration we set $N=10,000,D=3,J=3,T=2$. Since in this case the sufficient conditions for point identification are satisfied, any point from the argmin set $\hat{B}_{b}:=\arg\min_{\beta\in\mathbb{\mathbb{S}}^{D-1}}\hat{Q}_{b}\left(\beta\right)$ is a consistent estimator of $\beta_{0}$ for each round of simulation $b=1,..,M$. Specifically, we define
@@
 mid bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\beta_{0,d}\right)$0008 .0003 -.0067
 upper bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{u}-\beta_{0,d}\right)$0083 .0085 .0046
 lower bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{l}-\beta_{0,d}\right)$.0068 -.0080 -.0180
 mean(u--l) $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{u}-\hat{\beta}_{b,d}^{l}\right)$0152 .0165 .0226
standard deviation $\sqrt{\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\overline{\hat{\beta}_{d}^{m}}\right)^{2}}$0302 .0311 .0447
root MSE (by coordinate) $\left(\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\beta_{0,d}\right)^{2}\right)^{1/2}$0293 .0300 .0438
root MSE (whole vector) $\left(\frac{1}{M}\sum_{b=1}^{M}\left\|\hat{\beta}_{b}^{m}-\beta_{0}\right\|^{2}\right)^{1/2}$0606
mean normdeviations ($\begin{array}[]{c}\text{mean norm}\\ \text{deviations (MND) }\end{array}$ $\frac{1}{M}\sum_{b=1}^{M}\left\|\hat{\beta}_{b}^{m}-\beta_{0}\right\|$0525
Table 1 summarizes our baseline results. In the first row we use the middle point $\hat{\beta}^{m}$ along each dimension of $\hat{B}$ to calculate the average bias. The bias is very small across all three dimensions with a magnitude between -.0067 and .0008. The next two rows show the biases in estimating $\beta_{0,d}$ using $\hat{\beta}_{d}^{u}$ and $\hat{\beta}_{d}^{l}$ respectively and the biases are again close to zero. The fourth row measures the average width of the set estimator $\hat{B}$ along each dimension. It is relatively tight compared to the magnitude of $\beta_{0}$. The fifth and sixth row summarizes the standard deviation and root mean squared error for each coordinate of $\widehat{\beta}^{m}$. In the second part of Table 1 we report vector rMSE and MND based on $\hat{\beta}^{m}$, and our method performs very well.
Results Varying $N$
Next, we vary $N$ while maintaining $D=3,J=3,T=2$ to show how our method performs under different sample sizes. In addition to the baseline setup with $N=10,000$, we calculate mean absolute deviation (MAD), average size of the estimated set, rMSE and MND for $N=4,000$ and $N=1,000$. Results are summarized in Table 2.
Table 2: Performance under Varying $N$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N=10,000}$ 3.16 2.15 2.18 2.15
[…]
Table 2 provides numerical evidence that a larger $N$ helps with overall performance. The sum of absolute bias decreases from .0520 to .0077 when $N$ increases from $1,000$ to $10,000$. The average size of the estimated sets, rMSE, and MND follow a similar pattern. However, even with a relatively small $N=1,000$, our result is still informative and accurate, with the rMSE and MND being equal to .1320 and .1128, respectively. We note that here the total number of time periods $T$ is set to a minimum of 2. Our method can extract information from each of the $T\left(T-1\right)$ ordered pairs of time periods, thus a larger $T$ would in theory improve the performance of our estimators. See Appendix D for more simulation results with larger $T$.
Finally, we numerically investigate the speed of convergence when we increase $N$ from $1,000$ to $4,000$ and $10,000$ in the second part of Table 2. Compared with the case of $N_{0}=1,000$, the relative ratios of rMSE are 1.73 for $N=4,000$ and 2.18 for $N=10,000$, both of which lie between $\left(N/N_{0}\right)^{1/3}$ and $\left(N/N_{0}\right)^{1/2}$. A similar pattern is also found for calculations based on MND. These results indicate that our estimator achieves a convergence rate slower than the $N^{-1/2}$ but faster than the $N^{-1/3}$ rate.
@@ 6.1 Data and Methodology
We now provide an empirical application of the panel multinomial choice model and the estimation method we propose, using the Nielsen Retail Scanner Data on popcorn sales to explore the effects of display promotion effects. The data contains weekly information on store-level price, sales and display promotion status generated by about 35,000 participating retail stores with point-of-sale systems across the United States.
Among a huge variety of products we choose to focus on popcorn for two reasons. First, purchases of popcorn are more likely to be driven by temporary urges of consumption without too much dynamic planning. Second, there is good variation in the display promotion status of popcorn, which enables us to estimate how important special in-store displays affect consumer’s purchase decisions.
@@
DMA-level Market Share $s_{ijt}$ $25.00\%$ $21.59\%$ $0.07\%$ $96.69\%$
Priceijt .4924 .1803 .1094 1.3587
$\text{Promo}_{ijt}$0282 .0377 .$0$5000
$\text{Price}{}_{ijt}$$\times$ $\text{Promo}_{ijt}$0136 .0203 .$0$4505
To describe the methodology, we use the observed DMA-level market shares as an estimate of $s_{ijt}=\mathbb{E}\left[\left.y_{ijt}\right|{\bf X}_{it},{\bf A}_{i}\right].$ Under Assumption 3, we first estimate

### v3

@@ 4 General Econometric Framework of Multi-Index Single-Crossing Conditions
Similar to Example 2, define $\xi_{i}:=W_{i1}-Z_{i}^{{}^{\prime}}\gamma_{0}$ as the residual from the reduced-form linear projection of $W_{i1}$ on the instrumental variables $Z_{i}$, and assume that the endogeneity between $\epsilon_{i}$ and $W_{i1}$ is captured by the control function $\text{med}\left(\left.\epsilon_{i}\right|W_{i},Z_{i}\right)=\text{med}\left(\left.\epsilon_{i}\right|Z_{i},\xi_{i}\right)=\lambda\left(\alpha_{0}\xi_{i}\right)$, where $\lambda$ is an increasing function with normalization $\lambda\left(0\right)=0$. Writing $\overline{Z}_{i}:=\left(W_{i1},Z_{i}\right)$ and $\overline{\gamma}_{0}:=\left(\alpha_{0},-\alpha_{0}\gamma_{0}\right)^{{}^{\prime}}$, we have
$W_{i}^{{}^{\prime}}\beta_{0}>0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}>0\quad\Rightarrow\quad\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ $>0\text{ and}$
$W_{i}^{{}^{\prime}}\beta_{0}\leq 0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}\leq 0\quad\Rightarrow\quad\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ $=0,$
which can be viewed as a MISC condition with a weak $\leq$” side and $h_{0}\left(X_{i}\right):=\text{med}\left(\left.Y_{i}\right|W_{i},Z_{i}\right)$ given by the conditional median function.
Based on the MISC conditions (19), we can again obtain identifying restrictions by taking their logical contrapositions, which can be encoded algebraically in a similar way as in (1) and (14). Specifically, let $G$ be a one-sided sign-preserving function as in (14) and define
$\lambda\left(W_{i};\theta\right):=\prod_{j=1}^{J}\mathbf{\mathbbm{1}}\left\{W_{ij}^{{}^{\prime}}\theta_{j}\leq 0\right\}.$
@@
$h_{0}\left(X_{i}\right)<0$ $\ \Rightarrow\ \text{NOT}\ \left\{W_{ij}^{{}^{\prime}}\theta_{j}\geq 0\ \forall j\right\}.$
Furthermore, with $Q\left(\theta\right):=Q_{+}\left(\theta\right)+Q_{-}\left(\theta\right)$ where
[…]
we have $Q\left(\theta\right)\geq Q\left(\theta_{0}\right)=0$.
Proposition 2 generalizes Theorem 1. Notice that Proposition 2 applies to all functionals $h_{0}$ of the conditional distribution of $y_{i}$ given ${\bf X}_{i}$ that satisfy the MISC conditions.
One could also proceed with the two-step estimation procedure described in Section 3. Given a first-stage nonparametric estimator $\hat{h}$ of $h_{0}$, we can estimate $\theta_{0}$ (or the identified set) by minimizing the sample criterion $\hat{Q}\left(\theta\right):=\hat{Q}_{+}\left(\theta\right)+\hat{Q}_{-}\left(\theta\right)$ with
[…]
@@ 5 Simulation
We now switch back to the panel multinomial choice model introduced in Section 2 and examine the finite sample performance of our proposed estimator. For each DGP, we run $M=1,000$ simulations of model (1) with the following utility specification:
@@
in which $A_{i0}$ is an unobserved scale fixed effect that captures agent-level heteroskedasticity in utilities, and $A_{ij}$ is an unobserved location shifter specific to each agent-product pair. The ability to deal with nonlinear dependence caused by unobserved fixed effects in a relatively robust way is a distinctive feature of our method compared with existing approaches. To allow for such dependence, we generate correlation between the observable characteristics ${\bf X}_{i}$ and the fixed effects ${\bf A}_{i}$ via a latent variable $Z$. We draw $Z_{i}\sim\mathcal{N}$$\left(0,1\right)$ and let $A_{i2}=\left[Z_{i}\right]_{+}$. We construct $X_{ijt,2}=W_{ijt}+Z_{i}$ with $W_{ijt}\sim\mathcal{N}\left(0,2J\right)$. Thus, $X$ and $A$ are correlated via $Z$. The DGPs for the rest of $\bf A$} and $\bf X$} are: $A_{i0}\sim\mathcal{U}\left[2,2.5\right]$, $A_{i1}\equiv 0$, $A_{ij}\sim\mathcal{U}\left[-0.25,0.25\right]$ for $j\geq 3$, $X_{ijt,1}\sim\mathcal{U}\left[-1,1\right]$, $X_{ijt,d}\sim\mathcal{N}\left(0,1\right)$ for $d\geq 3$. Furthermore, we set $\overline{\beta}_{0}=\left(2,1,\ldots,1\right)^{{}^{\prime}}\in\mathbb{R}^{D}$ and $\beta_{0}=\overline{\beta}_{0}/\left\|\overline{\beta}_{0}\right\|$, and draw $\epsilon_{ijt}\sim TIEV\left(0,1\right)$. To summarize, for each of the $M=1,000$ simulations we first generate $\left(\beta_{0},{\bf X}_{it},{\bf A}_{i},\boldsymbol{\epsilon}_{it}\right)$ for all $(i,t)$ pairs. Then, we calculate the individual choice $\bf Y$} matrix according to model (1). Next, we compute $\hat{\beta}$ from the simulated observable data of $\left({\bf X},{\bf Y}\right)$. To obtain $\hat{\beta}$, we first use nonparametric regression with second-order polynomial basis functions with $\ell_{1}$-regularization and 10-fold cross validation to estimate $\gamma$. Then, we apply the adaptive-grid algorithm detailed in Section 3.3. Finally, we assess how well $\hat{\beta}$ performs compared with the true value $\beta_{0}$. [^17] In Appendix F, we provide additional simulation results. Specifically, we first present a graphical illustration of the identified set $B_{0}$ based on the population criterion (15). Second, we inspect how our estimator performs without point identification. Third, we vary $\left(D,J,T\right)$ to examine how robust our method is against various simulation specifications. Lastly, we include a simulation illustration of the robustness of our approach to the “Blue-Bus/Red-Bus” problem.
@@
For the baseline configuration, we set $N=10,000,\ D=3,\ J=3,\text{ and }T=2$. Since in this case the conditions for the point identification are satisfied, any point from the argmin set $\hat{B}_{b}:=\arg\min_{\beta\in\mathbb{\mathbb{S}}^{D-1}}\hat{Q}_{b}\left(\beta\right)$ is a consistent estimator of $\beta_{0}$ for each round of simulation $b=1,\ldots,M$. Specifically, we define
@@
 mid bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\beta_{0,d}\right)$ -0.0005 -0.0003 -0.0034
 upper bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{u}-\beta_{0,d}\right)$ 0.0075 0.0080 0.0083
 lower bias $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{l}-\beta_{0,d}\right)$ -0.0085 -0.0086 -0.0150
 mean(u-l) $\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{u}-\hat{\beta}_{b,d}^{l}\right)$ 0.0160 0.0166 0.0233
standard deviation $\sqrt{\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\overline{\hat{\beta}_{d}^{m}}\right)^{2}}$ 0.0299 0.0308 0.0431
root MSE (by coordinate) $\left(\frac{1}{M}\sum_{b=1}^{M}\left(\hat{\beta}_{b,d}^{m}-\beta_{0,d}\right)^{2}\right)^{1/2}$ 0.0288 0.0296 0.0417
root MSE (whole vector) $\left(\frac{1}{M}\sum_{b=1}^{M}\left\|\hat{\beta}_{b}^{m}-\beta_{0}\right\|^{2}\right)^{1/2}$ 0.0587
$\begin{array}[]{c}\text{mean norm}\\ \text{deviations (MND) }\end{array}$ $\frac{1}{M}\sum_{b=1}^{M}\left\|\hat{\beta}_{b}^{m}-\beta_{0}\right\|$ 0.0511
Table 1 summarizes our baseline results. In the first row we use the middle point $\hat{\beta}^{m}$ along each dimension of $\hat{B}$ to calculate the bias. The biases are very small across all three dimensions with a magnitude between -0.0034 and -0.0005. The next two rows show the biases in estimating $\beta_{0,d}$ using $\hat{\beta}_{d}^{u}$ and $\hat{\beta}_{d}^{l}$ respectively, which are again close to zero. The fourth row reports the average widths of the set $\hat{B}$ along each dimension. These widths are small relative to the magnitude of $\beta_{0}$. The fifth and sixth rows summarize the standard deviation and rMSE for each coordinate of $\widehat{\beta}^{m}$. In the second part of Table 1, we report the vector rMSE and MND based on $\hat{\beta}^{m}$, and the results suggest that our method performs well.
Results Varying NN
Next, we vary $N$ while maintaining D=3,J=3,and T=2D=3,\ J=3,\ $\text${and }T=2 to assess how our method performs under different sample sizes. In addition to the baseline setup with $N=10,000$, we calculate mean absolute deviation (MAD), average size of the estimated set, rMSE, and MND for $N=4,000$ and $N=1,000$. Results are summarized in Table 2.
Table 2: Performance under Varying NN
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N=10,000$} 3.16 2.15 2.33 2.27
N= 4, [^0] N=\ 4,000 2.00 1.59 1.84 1.78
Table 2 provides numerical evidence that a larger $N$ helps with overall performance. The sum of absolute bias decreases from 0.0606 to 0.0042 when $N$ increases from $1,000$ to $10,000$. The average size of the estimated sets, rMSE, and MND follow a similar pattern. Notably, even with a relatively small $N=1,000$, the results remain informative and reasonably accurate, with the rMSE and MND equal to 0.1369 and 0.1159, respectively. We note that $T$ is set to 2 here, which is the minimum required for our method to work. Since our method can extract information from each of the $T\left(T-1\right)$ ordered pairs of time periods, a larger $T$ would generally improve the performance of our estimators. Appendix F presents additional simulation results for a larger $T$.
Finally, we numerically investigate the speed of convergence when we increase $N$ from $1,000$ to $4,000$ and $10,000$ in the second part of Table 2. Compared with the case of $N_{0}=1,000$, the relative ratios of rMSE are 1.84 for $N=4,000$ and 2.33 for $N=10,000$, both of which lie between $\left(N/N_{0}\right)^{1/3}$ and $\left(N/N_{0}\right)^{1/2}$. A similar pattern is also observed for calculations based on MND. These results suggest that our estimator converges at a rate slower than $N^{-1/2}$ but faster than $N^{-1/3}$.
@@ 6.1 Data and Methodology
We now present an empirical application of the panel multinomial choice model and our proposed estimation method, using NielsenIQ Retail Scanner Data on popcorn sales to examine the effects of display promotions. The data contain weekly store-level information on prices, sales, and display promotion status, collected from approximately 35,000 participating retail stores with point-of-sale systems across the United States.
We focus on popcorn among the wide array of products for two reasons. First, popcorn purchases are more likely to be impulsive, with limited intertemporal planning. Second, popcorn exhibits substantial variation in in-store display promotions, allowing us to estimate how special displays influence consumers’ purchase decisions.
@@
DMA-level Market Share $s_{cjt}$ $25.06\%$ $21.59\%$ $0.08\%$ $96.69\%$
Pricecjt 0.4924 0.1803 0.1094 1.3587
$\text{Promo}_{cjt}$ 0.0282 0.0377 0.0000 0.5000
$\text{Price}{}_{cjt}$ $\times$ $\text{Promo}_{cjt}$ 0.0136 0.0203 0.0000 0.4505
Since the data are at the DMA level, while our approach was originally developed for individual-level data, we now describe how to adapt the method to the DMA-level setting. We treat the observed DMA-level market shares $s_{cjt}$ as noisy measurements [^19] Alternatively, with market-level data, one could treat the observed $s_{cjt}$ as a sufficiently good approximation of $\mathbb{E}\left[\left.y_{cjt}\right|{\bf X}_{ct},{\bf A}_{c}\right]$ as in Section 6.1 of Shi, Shum, and Song (2018), in which case our first-stage nonparametric regression is no longer required. We did not pursue this approach for two reasons: First, we do not wish to impose the assumption that the observed market shares are measured with negligible errors. Second, we intend this empirical exercise as an illustration of our two-stage procedure, and thus focus on a setting where the first-stage nonparametric regression is required. of $\mathbb{E}\left[\left.y_{cjt}\right|{\bf X}_{ct},{\bf A}_{c}\right]$, i.e.,
$s_{cjt}=\mathbb{E}\left[\left.y_{cjt}\right|{\bf X}_{ct},{\bf A}_{c}\right]+u_{cjt},\quad\text{with }\mathbb{E}\left[\left.u_{c}\right|{\bf X}_{c},{\bf A}_{c}\right]=0.$
Then, we use $s_{cjt}$ to nonparametrically estimate the following intertemporal difference:
$\mathbb{E}\left[\left.s_{cjt}-s_{cjs}\right|{\bf X}_{c,ts}\right]=\int\left(\mathbb{E}\left[\left.y_{cjt}\right|{\bf X}_{ct},{\bf A}_{c}\right]-\mathbb{E}\left[\left.y_{cjs}\right|{\bf X}_{cs},{\bf A}_{c}\right]\right)d\mathbb{P}\left(\left.{\bf A}_{c}\right|{\bf X}_{c,ts}\right).$
Specifically, we nonparametrically regress $\left(s_{cjt}-s_{cjs}\right)$ on the second-order polynomial basis functions of $\mathbf{X}_{c,ts}$ with $\ell_{1}$-regularization and 10-fold cross-validation to obtain an estimator $\hat{\gamma}_{j}$ of $\gamma_{j}\left(\overline{{\bf X}},\underline{{\bf X}}\right):=\mathbb{E}\left[\left.s_{cjt}-s_{cjs}\right|{\bf X}_{c,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]$. Finally, we plug $\hat{\gamma}$ into our second-stage algorithm and compute the (approximate) argmin set $\hat{B}_{\hat{c}}$.
6.2 Results and Discussion
We report our estimation results in Table 4. $\hat{\beta}_{\hat{c}}^{m}:=\frac{1}{2}\left(\hat{\beta}_{\hat{c}}^{l}+\hat{\beta}_{\hat{c}}^{u}\right)$ corresponds to the middle point of the (approximate) argmin set $\hat{B}_{\hat{c}}$ using our method. We show both the exact argmin set ($\hat{c}=0$) and the approximate argmin set with $\hat{c}=0.1\times N^{-\frac{1}{4}}\log\left(N\right)\approx 0.14$ for $N=205$. The estimated coefficients for Price (negative) and Promo (positive) are economically intuitive.
The most interesting result is the positive estimated coefficient on the interaction term $\text{Price}{}_{cjt}$ $\times$ $\text{Promo}_{cjt}$. An intuitive explanation for the positive sign is that by displaying certain products in front rows, consumers no longer see their price tags adjacent to those of their competitors, and thus become less price-sensitive for these specially promoted products.
Table 4: Empirical Illustration: Comparison of Results
$\hat{\beta}_{\hat{c}=0}^{m}$ $\hat{\beta}_{\hat{c}=0.14}^{m}$ $\hat{\beta}^{CyclicMono}$ $\hat{\beta}^{OLS}$ $\hat{\beta}^{OLS-FE}$ $\hat{\beta}^{MLogit-FE}$ $\hat{\beta}^{RCLM}$
Pricecjt -0.9351 -0.9283 -$\text{-}0.3781$ $0.0240$ -$\text{-}0.3807$ -$\text{-}0.6249$ -0.9705
$\text{Promo}_{cjt}$ 0.1793 0.1912 -$\text{-}0.0567$ $0.5760$ $0.5976$ $0.5881$ 0.2157
$\text{Price}{}_{cjt}$ $\times$ $\text{Promo}_{cjt}$ 0.2687 0.2505 $0.9240$ -$\text{-}0.8171$ -$\text{-}0.7057$ -$\text{-}0.5135$ 0.1078
Furthermore, we compare our $\hat{\beta}^{m}$ with the estimates obtained through four other methods, i.e., Cyclic Monotonicity (CM) based on Shi, Shum, and Song (2018) [^20] We use 2-week cycles for all available weeks in the data for the CM method., OLS, OLS with scalar-valued fixed effects (OLS-FE), the multinomial logit with fixed effects (MLogit-FE), and the random coefficients logit model (RCLM) [^21] See Appendix F.4 for the details of the RCLM estimator.. Results (normalized to $\mathbb{\mathbb{S}}^{D-1}$) are summarized in Table 4.